40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 18 22:49:06 2020, gautam, Update, ASC, Some thoughts about AS WFS electronics 
    Reply  Sun Nov 22 21:09:37 2020, gautam, Update, ASC, Planned mods for WFS head aLIGO_wfs_v5_40m.pdfTFs.pdfnoise.pdf
       Reply  Wed Nov 25 18:18:41 2020, gautam, Update, ASC, Planned mods for WFS head anodeVsOutput.pngsensitivity.png
          Reply  Sat Nov 28 21:44:53 2020, rana, Update, ASC, Planned mods for WFS head 
          Reply  Thu Dec 10 22:46:57 2020, gautam, Update, ASC, WFS head assembled aLIGO_wfs_v5_40m.pdfTF_meas.pdf
             Reply  Thu Dec 17 15:23:56 2020, gautam, Update, ASC, WFS head characterization oNoise.pdfZ_Hi.pdfZ_Low.pdfIMG_9030.jpgiNoise.pdf
    Reply  Wed Nov 25 18:30:23 2020, gautam, Update, ASC, Some thoughts about AS WFS electronics 
       Reply  Wed Dec 2 23:07:19 2020, gautam, Update, ASC, Electrical LO signal for AS WFS RF_Frequency_Source.pdf
          Reply  Thu Dec 3 11:02:35 2020, rana, Update, Electronics, Electrical LO signal for AS WFS 
          Reply  Thu Dec 3 13:55:52 2020, Koji, Update, ASC, Electrical LO signal for AS WFS 
             Reply  Thu Dec 3 20:38:46 2020, gautam, Update, ASC, Electrical LO signal for AS WFS RF_Frequency_Source.pdfdemodPath.pdfEOMpath.pdf
                Reply  Thu Dec 3 21:14:38 2020, Koji, Update, ASC, Electrical LO signal for AS WFS 
                   Reply  Thu Dec 3 21:44:49 2020, gautam, Update, ASC, Electrical LO signal for AS WFS 
                      Reply  Fri Dec 4 03:02:40 2020, Koji, Update, ASC, Electrical LO signal for AS WFS 
                         Reply  Fri Dec 4 22:41:56 2020, gautam, Update, ASC, Freq Gen Box revamp 
                            Reply  Sat Dec 5 20:44:35 2020, gautam, Update, ASC, Freq Gen Box re-installed 6x
                               Reply  Mon Dec 7 12:38:51 2020, gautam, Update, IOO, IMC loop char OLTFcomparison.pdfdemodMons.pdfOLTFcomparison.pdf
                               Reply  Mon Dec 7 22:54:30 2020, gautam, Update, LSC, Modulation depth measurement modDepth.pdf
                                  Reply  Sat Jan 16 18:59:44 2021, gautam, Update, LSC, Modulation depth measurement PLL.pdfmodDepth.pdfmodProfile.pdfEOMpath_postMod.pdf
Message ID: 15715     Entry time: Mon Dec 7 22:54:30 2020     In reply to: 15711     Reply to this: 15769
Author: gautam 
Type: Update 
Category: LSC 
Subject: Modulation depth measurement 

Summary:

I measured the modulation depth at 11 MHz andf 55 MHz using an optical beat + PLL setup. Both numbers are ~0.2 rad, which is consistent with previous numbers. More careful analysis forthcoming, but I think this supports my claim that the optical gain for the PDH locking loops should not have decreased.

Details:

  • For this measurement, I closed the PSL shutter between ~4pm and ~9pm local time. 
  • The photodiode used was the NF1611, which I assumed has a flat response in the 1-200 MHz band, and so did not apply any correction/calibration.
Attachment 1: modDepth.pdf  124 kB  | Hide | Hide all
modDepth.pdf
ELOG V3.1.3-